HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

GENERATING COMPLETE SETS OF BEYOND-VOIGT LINE-SHAPE

PARAMETERS FOR THE HITRAN DATABASE: DATASET STRUCTURE, TEMPERATURE-DEPENDENCIES, LINE-SHAPE MODEL

\underline{N. STOLARCZYK} , \underline{P. WCIS� O} , Institute of Physics, Faculty of Physics, Astronomy and Informatics, Nicolaus Copernicus University in Toruń, ul

Grudziadzka 5, 87-100 Toruń, Poland); L. S. ROTHMAN, \underline{R. HARGREAVES} , \underline{I. E. GORDON} , Harvard-Smithsonian Center for Astrophysics, Atomic and

Molecular Physics Division, Cambridge, MA, USA

The accurate characterization of collision-perturbed spectral lines from many molecular systems requires a profound understanding of the beyond-Voigt effects that influence the spectrum^{1}. These effects are described by eight essential spectral lineshape parameters: pressure broadening and shift (γ_{0} and δ_{0}), their speed dependencies (γ_{2} and δ_{2}), the real and imaginary parts of the complex Dicke parameter (ν̃_{opt}^{r} and ν̃_{opt}^{i}) and the complex line-mixing coefficient (y^{r} and y^{i}). Accurate knowledge of the temperature dependencies of these parameters is crucial for spectroscopic investigations of terrestrial^{2} and extraterrestrial^{3} atmospheres.

In my presentation, I will focus on the methodology employed to establish an extensive dataset of beyond-Voigt spectral line-shape parameters ^{4}. I will discuss interpolation and extrapolation schemes of the ab initio data to provide entries for the entire datasets in the HITRAN database^{5}. I will introduce the double-power-law (DPL) temperature-dependence model ^{6}, which offers a universal, straightforward, and accurate structure for each of the six line-shape parameters. I will discuss a new beyond-Voigt line-shape profile recommended for the HITRAN database.

1doi:10.1016/j.jqsrt.2016.01.024, P. Wcis�o, I.E. Gordon, H. Tran, Y. Tan, S.-M. Hu, A. Campargue, S. Kassi, D. Romanini, C. Hill, R.V. Kochanov and L.S. Rothman, JQSRT, 177, 75-91, (2016).

2doi:10.1016/j.crhy.2005.09.005, C.E. Miller, L.R. Brown, R.A. Toth, D.C. Benner, V. M. Devi, C R Phys, 6, 876-887, (2005).

3doi:10.1016/0019-1035(89)90062-6, W.M.H. Smith, C.P. Conner, J. Simon, W.V. Schempp, W.

Macy, Icarus, 81, 429-440, (1989).

4doi:10.1016/j.jqsrt.2020.107477, P. Wcis�o, F. Thibault, N. Stolarczyk, H. Jóźwiak, M. S�owiński, M. Gancewski, K. Stankiewicz, M. Konefa�, S. Kassi, A. Campargue, Y. Tan, J. Wang, K. Patkowski, R. Ciury�o, D. Lisak, R. Kochanov, L.S. Rothman, I.E. Gordon, JQSRT, 260, 107477, (2021).

5doi:10.1016/j.jqsrt.2021.107949, I.E. Gordon et al, JQSRT, 277, 107949, (2022).

6doi:10.1016/j.jqsrt.2019.106676, N. Stolarczyk, F. Thibault, H. Cybulski, H. Jóźwiak, G. Kowzan, B. Vispoel, I.E. Gordon, L.S. Rothman, R.R. Gamache, P. Wcis�o, JQSRT, 240, 106676, (2020).

p-number: p130  Submitted on Wed, 14 Jun 2023 14:17:28 +0200